HUBBLE SPACE TELESCOPE DAILY REPORT #919 PERIOD COVERED: 0800 EST Daily Status Report as of 181/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of PC 4714 (WF/PC cycle 2 Calibration: PC Internal Delta Flats) The Faint Object Camera was used to make five internal flat calibration images which will be used for the restoration of Earth flat images, these were altered due to changes in quantum efficiency which occurred after decontamination operations. The observations was received and displayed in the OSS. It was reported to look okay. 1.2 Completed FOC/96 4845 (Study of Optical Emission Associated With Radio Jets and Hot Spots) The Faint Object Camera was used to make four broad-band observations of the known radio source IC-5063 to study its current morphology, and to look for an optical counterpart to the radio jet. The observations were executed in fine lock. No anomalies were noted. 1.3 Completed WFC 4849 (GTO High Latitude Parallel Program Cycle 3 High) The Wide Field Camera was used to make four high latitude sky survey observations taken in parallel with the prime science observations. The observations were executed as planned in fine lock, and no anomalies were reported. 1.4 Completed PC 4809 (Cycle 2: WF/PC PSF Calibration Proposal) The Planetary Camera was used to make point spread function calibration image of standard star HD151406. The OSS monitored the observations and assessed the data quality. No problems were reported. 1.5 Completed FOC/96 4804 (UV Imaging of Nearby Galaxies with FOC/96) The Faint Object Camera was used to complete two sets of snapshot ultraviolet observations fo the nearby galaxy UGC-8 and UGC-786. The exposures lasted for ten minutes and used filter F220W. Both sets of observations were reported to be fine. 1.6 Completed WFC 4277 (Calibration of Nearby Galaxies Type IA Supernovae as Standard Candles) The Wide Field Camera was used to make a series of observations of Cepheid variable stars located inside fo galaxy NGC-5253. This is to help determine the accurate distance scale to this galaxy. The distance will be used to find the true luminosity of the type - Ia supernova which has been observed in this galaxy. The observations were executed as scheduled, and no problems were reported 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 6 Successful : 6 Scheduled Reacquisitions : 8 Successful : 8 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 9 Successful : 8 Failed FHST 1 & 3 Update, failed to perform a delay mode role update. 181/0445Z. 2.3 Significant HSTARs: None 2.4 Operations Note: Internal Simulation #4 (EVA 3) is currently under way today (Wednesday, June 30, 1993), and Simulation #10 (EVA 4) will be conducted tomorrow (Thursday, July 1, 1993). All operations are going as expected. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete FOC/96 4845 (Study of Optical Emissions Associated With Radio Jets and Hot Spots) 3.2 Complete WFC 4849 (GTO High Latitude Parallel Program, Cycle 3 High) 3.3 Uplink Updated Ephemeris Tables, Per MOSES 2871 3.4 Complete WFC 4713 (WF/PC Cycle 2 Calibration: WFC Internal Delta Flats) 3.5 Complete GHRS/2 4226 (Ultraviolet Spectroscopy of Hydrogen Poor Planetary Nebulae) 3.6 Complete WFC 4713 (WF/PC Cycle 2 Calibration: WFC Internal Delta Flats) 3.7 Complete GHRS/2 4510 (The Evolution of Boron: Implications for Galactic Evolution and Cosmos) 3.8 Complete PC 3603 (Snapshot of Protoplanetary Nebulae) 3.9 Complete PC/FOS 5021 (HST Observations of Comet Shoemaker-Levy (1993E)) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END- HUBBLE SPACE TELESCOPE DAILY REPORT #920 PERIOD COVERED: 0800 EST Daily Status Report as of 182/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed FOC/96 4845 (Study of Optical Emissions Associated With Radio Jets and Hot Spots) The Faint Object Camera was used to make three broad-band photometry observations of the known radio source NGC-5506 to study its current morphology, and to look for an optical counterpart to the radio jet and hot spots. The observations were executed in fine lock with no apparent problems. The observations were received and displayed in the OSS. It was reported that they could not tell if the intended target was in the aperture. 1.2 Completed WFC 4849 (GTO High Latitude Parallel Program, Cycle 3 High) The Wide Field Camera was used to make two high latitude sky survey observations taken in parallel with the prime science observations. The observations were executed in fine lock as planned. The observations were received and displayed in the OSS. They were reported to look okay. 1.3 Uplinked Updated Ephemeris Tables, Per MOSES 2871 1.4 Completed Two Sets of WFC 4713 (WF/PC Cycle 2 Calibration: WFC Internal Delta Flats) The Wide Field Camera was used to make a total of six internal flat calibrations images. All six images were reported to be normal. 1.5 Completed GHRS/2 4226 (Ultraviolet Spectroscopy of Hydrogen Poor Planetary Nebulae) The High Resolution spectrograph was used to make ultraviolet spectral observations of several emission knots and the central stars of the hydrogen-poor planetary nebula PK81-14D1. All observations were received in real time and were reported to look okay. 1.6 Completed GHRS/2 4510 (The Evolution of Boron: Implications for Galactic Evolution and Cosmos) The High Resolution Spectrograph was used to make six spectrographic observations of the late type star HD194598 to help study the chemical and dynamical evolution of the galaxy. The first five observations were reported to look okay. The peakup in the Small Science Aperture was successful. There was at least one recentering events which occurred during the last science observation. 1.7 Completed PC 3603 (Snapshot of Protoplanetary Nebulae) The Planetary Camera was used to make a snapshot image of the protoplanetary system PK074+02D1 to study its current morphology. No problems were reported. The image was received in the OSS and showed a nice planetary nebula. 1.8 Completed PC/FOS 5021 (HST Observations of Comet Shoemaker-Levy (1993E)) Both the Planetary Camera and the Faint Object Spectrograph were used to make observations of the Comet Shoemaker-Levy (aka, Comet 1993E). This comet broke into many fragments during a close approach to the planet Jupiter in 1992. The fragments are currently strung out in a line about one arcminute in length. The string of fragments was imaged with the planetary Camera, and spectra was obtained with the FOC Red Side detector. Of particular interest are the OH lines at 3085 A0, and possible 2825 A0 and 3144 A0. The principal investigator indicated that the initial pointing adjustment was not needed. The first three images were reported to look great! After examining the images a pointing correction was then uplinked to the spacecraft. No problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 6 Successful : 6 Scheduled Reacquisitions : 8 Successful : 8 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 10 Successful : 10 2.3 Significant HSTARs: None 2.4 Operations Note: None 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete FOC/96 4842 (Star formation Rates in Distant Cluster Galaxies - Cycle 3 High) 3.2 Complete WFC 4849 (GTO High Latitude Parallel Program Cycle 3 High) 3.3 Complete GHRS/2 4359 (Probing the Hot Interstellar Medium of the Largest LMC Sugergiants Shells) 3.4 Complete PC 3603 (Snapshot fo Protoplanetary Nebulae) 3.5 Complete WFC 3857 (Morphology of Galaxies in Clusters at Z = 0.5 - Continuation of 2373) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-